please help me solve this problem: the product of two consecutive positive integers is 29 more than their sum. find the integers.
Answer by mananth(16949)

let the consequtive integers bex & x+1
their product = x(x+1)
x(x+1)=x+x+1+29
x^2+x=2x+30
-2x
x^2-x=30
x^2-x-30=0
x^2-6x+5x-30=0
x(x-6)+5(x-6)=0
(x-6)(x+5)=0
x= 6 OR -5
6,7 & -5,-4 are the integers
m.ananth@hotmail.ca